18 Term with x 5y18 in the expansion of (-3x +y) b. Constant term in the expansion of- z 15
Answer #1

(a) Here, we will use Binomial Expansion in order for finding term with 5, 13 .

Binomial Expansion Formula is as : a +b) = (Pao

So, let us get started for (-3r -y)18

impliesじ ) 0 +   inom{18}{1}(-3x)^{17}(rac{1}{4}y)^1 + 23)16 1 + inom{18}{3}(-3x)^{15}(rac{1}{4}y)^3 + ... + 3 5 + inom{18}{14}(-3x)^{4}(rac{1}{4}y)^{14} + ... + inom{18}{18}(-3x)^{0}(rac{1}{4}y)^{18}   

Here, in above expansion we can see that 14th term having 5, 13 ​​​​​​ term i.e. 3 5 =   8568(-3x)^{5}(rac{1}{4}y)^{13} ,

Thus , the 14th term having 5, 13 term .

(b) Here , we are given   (rac{2}{z^{2}}- z)^{15} ,

we will expand this using Binomial expansion and

Let us consider the General Term for above Binomial is :

T_{r+1} =  inom{15}{r}(rac{2}{z^{2}})^{15 - r}(-z)^{r}  

T_{r+1} = inom{15}{r}(2)^{15-r}(-z)^{r}(z)^{2r-30}   

T_{r+1} = inom{15}{r}(2)^{15-r}(-z)^{3r-30}   

Now , for a term in the expansion to be constant , the power of Z should be 0 .

So, 3r - 30 = 0 enter implies r = 10 .

Therefore , the 11th term is the constant term in the expansion .

T_{10+1} = T_{11}

implies T_{11} = inom{15}{10}(2)^{15-10} (-z)^{3*10 - 30}

implies T_{11} = inom{15}{10}(2)^{15-10} (z)^{0}(-1)

implies T_{11} = -3003